### v1

@@ Abstract
An enhancement in seismic measuring instrumentation has been proven to have implications in the quantity of observed earthquakes, since denser networks usually allow recording more events. However, phenomena such as strong earthquakes or even aseismic transients, as slow slip earthquakes, may alter the occurrence of earthquakes. In the field of seismology, it is a standard practice to model background seismicity as a Poisson process. Based on this idea, this work proposes a model that can incorporate the evolving spatial intensity of Poisson processes over time (i.e., we include temporal changes in the background seismicity when modeling). In recent years, novel methodologies have been developed for quantifying the uncertainty in the estimation of the background seismicity in homogeneous cases using Bayesian non-parametric techniques. This work proposes a novel methodology based on graphical Dirichlet processes for incorporating spatial and temporal inhomogeneities in background seismicity. The proposed model in this work is applied to study the seismicity in the southern Mexico, using recorded data from 2000 to 2015.
@@ 1 Introduction
Since its introduction, the Epidemic Type Aftershock Sequence (ETAS) model [ogata1988statistical] has proved to be one of the most important tools for assessing stochastic seismicity, in which earthquake activity is modeled using Hawkes’ process [Hawkes1971Spectra]. In this context, the work by [ogata1988statistical] assumed that the mainshocks (or background earthquakes) originate from a Poisson process.
Recently, due to the importance of having uncertainty quantification of the estimators, such authors as [ross2022semiparametric] and [molkenthin2022gp] have worked with Bayesian non-parametric approaches to model a temporally homogeneous background intensity function $\mu(x,y)$ in equation (1) using Dirichlet processes mixtures (DPM) [antoniak1974mixtures] and Gaussian processes, respectively.
Since the homogeneity of the ETAS model can be affected by tectonic or pore pressure changes [hainzl2005detecting, hainzl2013impact, marsan2013monitoring], such phenomena may have implications for the occurrence of tremors. However, these changes can be produced by strong earthquakes [cruz2021short], and can also lead to an increase in the seismicity rate [nishikawa2017detection].
Furthermore, the presence of slow slip events—slow fault slip phenomena that mainly occur along the plate interface—can alter seismic activity [marsan2013slow, fukuda2018variability, nishikawa2023development], as well as periods of microearthquake quiescence [matthews1988statistical]. Moreover, not only variations in seismicity can affect ETAS model estimates, but also factors such as short-term aftershock incompleteness [hainzl2016apparent, asayesh2025improved] and changes in the cut-off magnitude [hainzl2022etas]. These temporary changes in activity emphasize the need for non-homogeneous models over time for the background activity function, such as the one introduced in this paper.
@@
  • We propose a nonparametric Bayesian approach to estimate the background seismic function based on Graphical Dirichlet Processes (GDP) [chakrabarti2024graphical], which allows for inhomogeneities in both space and time.
@@
As discussed in detail in Section 2, the usage of hidden variables in our methodology can be extended and incorporated in the context of the ETAS model, enhancing the methodology by [ross2022semiparametric].
In the frequentist context, [li2020space] have used piecewise constant functions and kernel density estimators to model an inhomogeneous space-time background activity function in the context of the Hawkes process. They assumed that
@@
In the non-parametric Bayesian framework, the estimation of $\mu(x,y,t)$ can be done using the Generalized Spatial Dirichlet Process introduced by [duan2007generalized]. Their approach allows modeling different probability density functions across space and discretized time, where the weights and point masses in the stick-breaking representation [sethuraman1994constructive] are assumed to vary smoothly over space and time. However, their formulation requires modeling the weights through an auxiliary thresholding scheme based on a collection of independent stationary Gaussian random fields. Since we keep the mass points fixed in our work, we have chosen to use the representation proposed by [chakrabarti2024graphical], which allows the weights to vary through a hierarchical model with beta distributions.
It is also worth mentioning the [fonseca2017dynamic] work, which deals with the spatio-temporal problem that occurs when point data are not directly observed and only aggregated counts are available. In this context, [kottas2007bayesian] proposed a non-parametric Bayesian approach to estimate the time-homogeneous spatial Poisson-process intensity function using Dirichlet processes, work that has been extended to the non-homogeneous case by [xiao2015modeling] using dependent Dirichlet processes (DDP) [maceachern1999dependent]. If the reader is interested in DDP and their stick-breaking representation, we suggest consulting the work [quintana2022dependent].
The main difference between the methodology by [xiao2015modeling] and the one proposed in our work lies on the DDP structure; while Kottas keeps the weights of the mixing elements fixed over time during the Dirichlet process (while varying their locations), we keep the locations fixed and move the weights, which will be discussed in detail in section 2.
@@ 2 Methodology
As is pointed out by [Fox2016Spatially], the usual approach to model the mainshocks (i.e., earthquakes that are not produced by other earthquakes, as the aftershocks), is to use an inhomogeneous Poisson Process in space and constant over time with intensity $\mu(x,y)$, and the rate of aftershocks is modeled using a triggering function $\nu$ depending on the magnitude ($M$) which decays according to the spatial and temporal distance between earthquakes. Then, the intensity function in the ETAS model is given by
$\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})=\mu(x,y)+\sum_{\{i:t_{i}<t\}}\nu(t-t_{i},x-x_{i},y-y_{i},M_{i})$ (1)
where $(x_{i},y_{i})$ are the coordinates of the epicenter and $t_{i}$ the time of occurrence of the event $i$, with i=1,2,,Ni=1,2,...,N. Physically, $\mu(x,y)$ denotes the background seismicity, which refers to the earthquake rate that is unaffected by the occurrence of any other seismic event. The triggering function $\nu$ is described as the decay of an earthquake’s ability to generate aftershocks over time (i.e., it is based on the Omori Law [utsu1995centenary]). Then, only earthquakes with magnitude exceeding an umbral $M_{u}$ are considered, based on the magnitude completeness provided by the Gutenberg-Richter Law[gutenberg1949seismicity]. Then, the log-likelihood of the model is given by
$\ell(\mu,\nu):=\sum_{k=1}^{n}\log(\lambda^{*}_{\mu,\nu}(t_{k},x_{k},y_{k}|H_{t_{k}}))-\int_{0}^{T}\iint_{S}\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})dxdydt,$ (2)
As we mentioned in Section 1, the catalog from [sawires2019updated] allows us to focus solely on the Poissonian intensity function. Therefore, in the rest of the text, we will work with a Non-homogeneous Poisson Process (NHPP) with intensity function defined as:
@@
to model the background seismic function.
It is important to note that if we have a sample $X=\{(x_{i},y_{i},t_{i})\}$, where $(x_{i},y_{i})$ are observed in a region $D\subset\mathbb{R}^{2}$, with i=1,2,,Ni=1,2,...,N and in a time interval $(0,T)$ from a NHPP the likelihood [daley2003introduction] is given by
$L(X|\mu)=e^{\iint_{D}\int_{0}^{T}\mu(x,y,t)dtdxdy}\prod_{i=1}^{N}\mu(x_{i},y_{i},t_{i}),$ (4)
which, in general, is enormously computationally expensive to compute because it requires solving a triple integral. Consequently, authors such as [kottas2007bayesian, zhuang2002stochastic, ross2022semiparametric], among others, have assumed, in the stationary case, that $\mu(x,y)$ is a bounded function. This form allows modeling it as
$\mu(x,y)=\gamma f(x,y)$
@@
As we discussed in 1 following the idea of writing $\mu(x,y)=\gamma f(x,y)$, authors as [li2024multivariate] and [xiao2015modeling] have proposed defining $\mu$ piecewise constant in time, i.e., if we have a partition of the interval (0,T) given by times $0=S_{1},S_{2},...,S_{p+1}=T$, then we can define $\mu$ as
$\mu(x,y,t)=\sum_{p=1}^{P}\gamma_{p}f_{p}(x,y)\mathbbm{1}(S_{p}<t<S_{p+1})$ (5)
where $\mathbbm{1}(\cdot)$ is the indicator function, which is the definition of $\mu$ that is used in our work. This definition aims to work with the inhomogeneities over space and time, but also allows us to rewrite the equation (4) as
@@
where $I_{p}=(S_{p},S_{p+1})$, $|I_{p}|$ is its longitude, $X_{p}=\{(x_{p,i},y_{p,i})\}$ with $i=1,...,n_{p}$ is the subsample of $X$ of elements where $t_{i}$ is in $I_{p}$ for p=1,,Pp=1,...,P.
As can be seen in this definition of $\mu$, we have created non-exchangeable subsamples of $X$ that are dependent on $\prod_{i=1}^{n_{p}}f_{p}(x_{p,i},y_{p,i})$, which is a reasonable model if dependency is assumed. The temporal dependency used in this work can be described as a particular case of GDP, which directed acyclic graph is shown in Figure 1.
The GDP [chakrabarti2024graphical] allows working with a more general DAG. Nevertheless, for our purposes, this version is adequate, and it also coincides with structures previously studied in the nested Chinese restaurant process [blei2010nested] or the Nested Hierarchical Dirichlet Processes [paisley2014nested] that are englobed in the GDP framework.
@@
Figure 1: Graphical Dirichlet Process denoting time dependency.
As mentioned by [muller2015bayesian], and since the Dirichlet processes are discrete with probability one, it is awkward to estimate continuous densities. This trouble could be fixed using DPM, which is generalized in [chakrabarti2024graphical] to the GDP. They proved that the structure presented in figure 1 used to define a GDP mixture is
$G_{1}|\alpha_{1},G_{0}$ $\sim DP(\alpha_{1},G_{0})$
$G_{p}|\alpha_{p},G_{p-1}$ $\sim DP(\alpha_{p},G_{p-1})$
@@
$\theta_{p,l}|G_{p}$ $\stackrel{{\scriptstyle\text{ind}}}{{\sim}}G_{p}$
$(x_{p,i},y_{p,i})|\{\theta_{p,l}\}_{l=1}^{L}$ $\sim F(\{\theta_{p,l}\}_{l=1}^{L})$
which can be recovered as the limit of a hierarchical finite model when the number of $L$ mixture elements tends to infinity. Then, the addition of $\gamma_{p}$ variables, a priori independent to all $G_{p}$ processes in our hierarchical model, is given by
$\alpha_{1}|\alpha_{0}$ $\sim\text{Gamma}(\alpha_{0},1)$ (7)
$\beta_{1}|\alpha_{1}$ $\sim\text{Dir}\Big(\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L}\Big)$
$\alpha_{t}|\alpha_{t-1}$ $\sim\text{Gamma}(\alpha_{t-1},1)$
$\beta_{p}|\alpha_{p},\beta_{p-1}$ $\sim\text{Dir}(\alpha_{p}(\beta_{p-1,1},\beta_{p-1,2},...,\beta_{p-1,L}))$
$\psi_{l}|G_{0}$ $\sim G_{0}$
$z_{pi}|\beta_{t}$ $\sim\text{Cat}(1:L,\beta_{p})$
$\gamma_{p}$ $\sim\text{Gamma}(\gamma_{0}k,k)$
$\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{p}}$ $\sim\mathcal{F}(z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p})\text{ for all $p$ in \{1,2,..,P\}}$
where $\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{t}}\sim\mathcal{F}(z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{t})$ has the following probability density function
$f_{x}((x_{p,1},y_{p,1}),...,(x_{p,n_{p}},y_{p,n_{p}}))=e^{-\gamma_{p}|I_{p}|}\gamma_{t}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big).$
@@
The main reason for assuming that $f_{\psi_{l}}$ is Gaussian is to obtain the full conditional of $\psi_{l}$ using conjugate analysis, which will allow its incorporation into a Markov Chain Monte Carlo (MCMC) straightforwardly.
@@
For all our examples, we define $G_{0}\sim\text{NIW}(\mu_{0},\eta,\Sigma_{0},\nu)$ where NIW denotes a Normal Inverse Wishart distribution, also we denote his density by $f_{\text{NIW}}(\cdot)$. Therefore, our posterior distribution is given by
$f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|\boldsymbol{Y})$ $\propto\alpha_{1}^{\alpha_{0}-1}e^{-\alpha_{1}}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}-1}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})}\Big(\prod_{p=1}^{P-1}\frac{\alpha_{p+1}^{\alpha_{p}-1}e^{-\alpha_{p+1}}}{\Gamma(\alpha_{p})}\Big)$
@@
$\prod_{p=1}^{P}e^{-\gamma_{p}|I_{p}|}\gamma_{p}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)$

### v2

@@ Abstract
Enhancements in seismic measuring instrumentation have been proved to have implications in the quantity of observed earthquakes, since denser networks usually allow recording more events. However, phenomena such as strong earthquakes or even aseismic transients, as slow slip earthquakes, may alter the occurrence of earthquakes. In the field of seismology, it is a standard practice to model background seismicity as a Poisson process. Based on this idea, this work proposes a model that can incorporate the evolving spatial intensity of Poisson processes over time (i.e., we include temporal changes in the background seismicity when modeling). In recent years, novel methodologies have been developed for quantifying the uncertainty in the estimation of the background seismicity in homogeneous cases using Bayesian nonparametric techniques. This work proposes a novel methodology based on graphical Dirichlet processes for incorporating spatial and temporal inhomogeneities in background seismicity. The proposed model in this work is applied to study the seismicity in southern Mexico, using recorded data from 2000 to 2015.
@@ 1 Introduction
Since its introduction, the Epidemic Type Aftershock Sequence (ETAS) model [ogata1988statistical] has proved to be one of the most important tools for assessing stochastic seismicity, in which earthquake activity is modeled using Hawkes process [Hawkes1971Spectra]. In this context, the work by [ogata1988statistical] assumed that the mainshocks (or background earthquakes) originate from a Poisson process.
Recently, due to the importance of having uncertainty quantification of the estimators, authors such as [ross2022semiparametric] and [molkenthin2022gp] have worked with Bayesian nonparametric approaches to model a temporally homogeneous background intensity function $\mu(x,y)$ in equation (1) using Dirichlet processes mixtures (DPM) [antoniak1974mixtures] and Gaussian processes, respectively.
Since the homogeneity of the ETAS model can be affected by tectonic or pore pressure changes [hainzl2005detecting, hainzl2013impact, marsan2013monitoring], such phenomena may have implications for the occurrence of tremors. However, these changes can be produced by strong earthquakes [cruz2021short], and can also lead to an increase in the seismicity rate [nishikawa2017detection, banales2025inhomogeneous].
Furthermore, the presence of slow slip events—slow fault slip phenomena that mainly occur along the plate interface—can alter seismic activity [marsan2013slow, fukuda2018variability, nishikawa2023development], as well as periods of microearthquake quiescence [matthews1988statistical]. Moreover, not only variations in seismicity can affect ETAS model estimates, but also factors such as short-term aftershock incompleteness [hainzl2016apparent, asayesh2025improved] and changes in the cut-off magnitude [hainzl2022etas]. These temporary changes in activity emphasize the need for inhomogeneous models over time for the background seismicity function, such as the one introduced in this paper.
@@
  1. 1.
    We propose a nonparametric Bayesian approach to estimate the background seismicity function based on Graphical Dirichlet Processes (GDP) [chakrabarti2024graphical], which allows for inhomogeneities in both space and time.
  2. 2.
@@
In the seismological literature, there are two main approaches to distinguish background events from aftershocks: those based on deterministic spatio-temporal filtering, as proposed in [gardner1974sequence] and used in [sawires2019updated], and those based on stochastic models following the ETAS model proposed in [ogata1988likelihood], as in [zhuang2002stochastic, schoenberg2013facilitated]. It is important to emphasize that earthquake modeling corresponds to a self-exciting point process, and while, as previously mentioned above in point 3, we will use the catalog of [sawires2019updated], any assumption about the background intensity function may lead to potential biases in the estimation of the triggering effects.
While the catalog of [sawires2019updated] allows us to focus only on the Poissonian part of the problem, in Section 2 we describe how our work can generate extensions of the model proposed in [ross2022semiparametric], based on the ETAS model, by incorporating hidden variables in a manner similar to that developed in [schoenberg2013facilitated, banales2025inhomogeneous].
In the frequentist context, [li2020space] has used piecewise constant functions and kernel density estimators to model an inhomogeneous space-time background activity function in the context of the Hawkes process. They assumed that
@@
In the nonparametric Bayesian framework, the estimation of $\mu(x,y,t)$ can be done using the Generalized Spatial Dirichlet Process introduced by [duan2007generalized]. Their approach allows modeling different probability density functions across space and discretized time, where the weights and point masses in the stick-breaking representation [sethuraman1994constructive] are assumed to vary smoothly over space and time. However, their formulation requires modeling the weights through an auxiliary threshold scheme based on a collection of independent stationary Gaussian random fields. Since we keep the mass points fixed in our work, we have chosen to use the representation proposed by [chakrabarti2024graphical], which allows the weights to vary through a hierarchical model with beta distributions.
It is also worth mentioning the work by [fonseca2017dynamic], which deals with the spatio-temporal problem that occurs when point data are not directly observed and only aggregated counts are available. In this context, [kottas2007bayesian] proposed a nonparametric Bayesian approach to estimate the time-homogeneous spatial Poisson-process intensity function using Dirichlet processes, work that has been extended to the inhomogeneous case by [xiao2015modeling] using dependent Dirichlet processes (DDP) [maceachern1999dependent]. If the reader is interested in DDP and their stick-breaking representation, we suggest consulting [quintana2022dependent].
The main difference between the methodology of [xiao2015modeling] and the one proposed here lies in the DDP structure; while [xiao2015modeling] keeps the mixture weights fixed over time and allows the component locations to vary, we instead keep the locations fixed and allow the weights to evolve over time, as discussed in Section 2.
@@
The modeling of spatio-temporal point processes is not limited to seismology. In particular, recent works such as [gilardi2024nonseparable, medialdea2025fully] have proposed models based on spatial Cox processes [Diggle2013LGCP] to model ambulance interventions and wildfires. Using a frequentist approach, their proposals also allow the incorporation of exogenous variables and planar geometries.
Also, in the context of self-exciting point processes, [briz2024point] proposed a purely temporal model that allows the self-exciting mechanism to be activated and deactivated over time, thereby reducing underestimation of the productivity parameter in a seismological context.
Finally, we mention the works of [zhuang2019semiparametric, escudero2025crime], which focused on modeling crime data, using log-Gaussian Cox and Hawkes processes with inhomogeneous separable spatio-temporal background intensities. This contrasts with the non-separable intensity proposed in this work based on a GDP, which highlights the novelty of our approach.
@@ 2 Methodology
As is pointed out by [Fox2016Spatially], the usual approach to model the mainshocks (i.e., earthquakes that are not triggered by other earthquakes, unlike aftershocks), is to use an inhomogeneous Poisson Process in space and constant over time with intensity $\mu(x,y)$, and the rate of aftershocks is modeled using a triggering function $\nu$ depending on the magnitude ($M$) which decays according to the spatial and temporal distance between earthquakes. Then, the intensity function in the ETAS model is given by
$\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})=\mu(x,y)+\sum_{\{i:t_{i}<t\}}\nu(t-t_{i},x-x_{i},y-y_{i},M_{i}),$ (1)
where $(x_{i},y_{i})$ are the coordinates of the epicenter, $t_{i}$ the time of occurrence of the event $i$, with i=1,2,,Ni=1,2,...,N, and $H_{t}$ denotes the history of earthquakes locations, times and magnitudes up to time $t$. Physically, $\mu(x,y)$ denotes the background seismicity, which refers to the earthquake rate that is unaffected by the occurrence of any other seismic event. The triggering function $\nu$ is described as the decay of an earthquake’s ability to generate aftershocks over time (i.e., it is based on the Omori Law [utsu1995centenary]). Then, only earthquakes with magnitude exceeding a $M_{u}$ threshold are considered, based on the magnitude completeness provided by the Gutenberg-Richter Law[gutenberg1949seismicity]. Then, the log-likelihood of the model is given by
$\ell(\mu,\nu):=\sum_{k=1}^{n}\log(\lambda^{*}_{\mu,\nu}(t_{k},x_{k},y_{k}|H_{t_{k}}))-\int_{0}^{T}\iint_{D}\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})dxdydt,$ (2)
where $(0,T)$ is the observation time interval and $D\subset\mathbb{R}^{2}$ denotes the observed region.
As we mentioned in Section 1, the catalog from [sawires2019updated] allows us to focus solely on the Poissonian intensity function. Therefore, in the rest of the text, we will work with a inhomogeneous Poisson Process (NHPP) with intensity function defined as:
@@
to model the background seismicity function.
It is important to note that if we have a sample $X=\{(x_{i},y_{i},t_{i})\}$, where $(x_{i},y_{i})$ are observed in a region $D\subset\mathbb{R}^{2}$, with i=1,2,,Ni=1,2,...,N and in a time interval $(0,T)$ from a NHPP, the likelihood [daley2003introduction] is given by
$L(X|\mu)=e^{-\int_{0}^{T}\iint_{D}\mu(x,y,t)dtdxdy}\prod_{i=1}^{N}\mu(x_{i},y_{i},t_{i}),$ (4)
which, in general, is enormously computationally expensive because it requires solving a triple integral. Consequently, authors such as [kottas2007bayesian, zhuang2002stochastic, ross2022semiparametric], among others, have assumed, in the stationary case, that $\mu(x,y)$ is a bounded function. This form allows modeling it as
$\mu(x,y)=\gamma f(x,y),$
@@
As discussed in Section 1 following the idea of writing $\mu(x,y)=\gamma f(x,y)$, authors such as [li2024multivariate] and [xiao2015modeling] have proposed defining $\mu$ piecewise constant in time, i.e., if we have a partition of the interval (0,T) given by times $0=S_{1},S_{2},...,S_{p+1}=T$, then we can define $\mu$ as
$\mu(x,y,t)=\sum_{p=1}^{P}\gamma_{p}f_{p}(x,y)\mathbbm{1}(S_{p}<t<S_{p+1}),$ (5)
where $\mathbbm{1}(\cdot)$ is the indicator function, which is the definition of $\mu$ that is used in our work. This definition aims to work with the inhomogeneities over space and time, but also allows us to rewrite Equation (4) as
@@
where $I_{p}=(S_{p},S_{p+1})$, $|I_{p}|$ is the width of the partition, $X_{p}=\{(x_{p,i},y_{p,i})\}$ with $i=1,...,n_{p}$ is the subsample of $X$ of elements where $t_{i}$ is in $I_{p}$ for p=1,,Pp=1,...,P.
As can be seen in this definition of $\mu$, we have created non-exchangeable subsamples of $X$ that are dependent on $\prod_{i=1}^{n_{p}}f_{p}(x_{p,i},y_{p,i})$, which is a reasonable model if dependency is assumed. The temporal dependency used in this work can be described as a particular case of GDP, for which the directed acyclic graph is shown in Figure 1.
The GDP [chakrabarti2024graphical] allows working with a more general DAG. Nevertheless, for our purposes, this version is adequate, and it also coincides with structures previously studied in the nested Chinese restaurant process [blei2010nested] or the nested Hierarchical Dirichlet Processes [paisley2014nested] that are encompassed in the GDP framework.
@@
Figure 1: Graphical Dirichlet Process denoting time dependency
As mentioned by [muller2015bayesian], and since Dirichlet processes are discrete with probability one, it is awkward to estimate continuous densities. This limitation can be overcome using DPM, which is generalized in [chakrabarti2024graphical] to the GDP. The authors of [chakrabarti2024graphical] proved that the structure presented in Figure 1 used to define a GDP mixture is
$G_{1}|\alpha_{1},G_{0}$ $\sim DP(\alpha_{1},G_{0}),$
$G_{p}|\alpha_{p},G_{p-1}$ $\sim DP(\alpha_{p},G_{p-1}),$
@@
$\theta_{p,i}|G_{p}$ $\stackrel{{\scriptstyle\text{ind}}}{{\sim}}G_{p},$
$(x_{p,i},y_{p,i})|\theta_{p,i}$ $\sim F(\theta_{p,i}),$
which, by the stick-breaking representation, can be recovered as the limit of the hierarchical finite mixture model presented in Equation (7) when the number of $L$ mixture elements tends to infinity. Then, the addition of $\gamma_{p}$ variables, a priori independent of all $G_{p}$ processes in our hierarchical model, is given by
$\alpha_{1}|\alpha_{0}$ $\sim\text{Gamma}(\alpha_{0},1),$ (7)
$\beta_{1}|\alpha_{1}$ $\sim\text{Dir}\Big(\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L}\Big),$
$\alpha_{t}|\alpha_{t-1}$ $\sim\text{Gamma}(\alpha_{t-1},1),$
$\beta_{p}|\alpha_{p},\beta_{p-1}$ $\sim\text{Dir}(\alpha_{p}(\beta_{p-1,1},\beta_{p-1,2},...,\beta_{p-1,L})),$
$\psi_{l}|G_{0}$ $\sim G_{0},$
$z_{p,i}|\beta_{p}$ $\sim\text{Cat}(1:L,\beta_{p}),$
$\gamma_{p}$ $\sim\text{Gamma}(\gamma_{0}k,k),$
$\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{p}}$ $\sim\mathcal{F}(\{z_{p,i}\}_{i=1}^{n_{p}},\{\psi_{l}\}_{l=1}^{L},\gamma_{p})\text{ for all $p$ in \{1,2,..,P\}},$
where $\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{p}}\sim\mathcal{F}(\{z_{p,i}\}_{i=1}^{n_{p}},\{\psi_{l}\}_{l=1}^{L},\gamma_{p})$ has the following probability density function
$f((x_{p,1},y_{p,1}),...,(x_{p,n_{p}},y_{p,n_{p}}))=e^{-\gamma_{p}|I_{p}|}\gamma_{t}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big),$
and as we previously introduced in Equation (6) we have
$f(X|\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z})=\prod_{p=1}^{P}\Big\{e^{-\gamma_{p}|I_{p}|}\gamma_{t}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)\Big\}.$
@@
The main reason for assuming that $f_{\psi_{l}}$ is Gaussian is to obtain the full conditional of $\psi_{l}$ using conjugate analysis, which will allow its straightforward incorporation into a Markov chain Monte Carlo (MCMC) scheme.
@@
For all our examples, we define $G_{0}\sim\text{NIW}(\mu_{0},\eta,\Sigma_{0},\nu)$ where NIW denotes a Normal Inverse Wishart distribution, we also denote this density by $f_{\text{NIW}}(\cdot)$. While the prior distribution reflects our uncertainty about the parameters before observing the data, this uncertainty is updated once the sample is observed through Bayes’ theorem, as detailed in [bernardo1994bayesian], which defines the posterior distribution, i.e.,
$f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|X)=\frac{f(X|\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z})f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z})}{f(X)}.$
Since the denominator is only a normalizing constant and does not depend on the parameters of interest, the posterior distribution is fully characterized by the following expression
$f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|X)$ $\propto f(X|\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z})f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z})$ (8)
$=\alpha_{1}^{\alpha_{0}-1}e^{-\alpha_{1}}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}-1}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})}\Big(\prod_{p=1}^{P-1}\frac{\alpha_{p+1}^{\alpha_{p}-1}e^{-\alpha_{p+1}}}{\Gamma(\alpha_{p})}\Big)$
@@
$\prod_{p=1}^{P}e^{-\gamma_{p}|I_{p}|}\gamma_{p}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big),$